0.3(4+x)=4.5 could you show it step by step
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

0.3(4+x)=4.5
1.2+0.3x=4.5
0.3x=3.3
x=11

Answer by stanbon(75887)

0.3(4+x)=4.5
---
Divide both sides by 0.3 to get:
(4+x) = 15
Subtract 4 from both sides to get:
x = 11
==============
Cheers,
Stan H.